### v1

# FLARE v2: A Recursive Framework for Program Comprehension Across Languages and Levels of Abstraction

Abstract
Building on the classroom framework reported in heath2025flare, this paper proposes FLARE v2 as a recursive, semiotically informed account of how program meaning is constructed. It reinterprets the descriptive tiers of FLARE v1 as instances of a single generative operation: identify elements (characterised by the four properties Receives, Sends, Effects, Shares); analyse their bindings along two dimensions (Causal-Temporal and Communicative); and recognise the new element that emerges. The Causal-Temporal dimension encompasses three subtypes - Sequential, Branch, and Event - that together account for control flow in both procedural and event-driven environments. A Compositional Ladder provides a visual parallel between literacy progressions and programming structures, illustrating how recursive composition operates from blocks and statements through segments, systems, and services. The framework aims to address conceptual and cognitive-load limitations reported in FLARE v1 and is situated within semiotic and program-comprehension theory. FLARE v2 is presented as a conceptual lens with potential implications for pedagogy and curriculum design; implementation and empirical evaluation are left for future work.
1 Introduction
Teaching programming requires teachers to help pupils make sense of program behaviour, yet many computing teachers are non-specialists. The Royal Society reported that many computing teachers felt under-supported and lacked specialist background in computer science (royalsociety2017), and a recent follow-up review found that only around a third of England’s computer science ITT recruitment target is being met and that just 40% of computing teachers hold a relevant post-A-level qualification (royalsociety2025). In addition, much of a program’s runtime behaviour is not directly visible in the source text. Sorva argues that, without explicit teaching about a notional machine, novices form incomplete and unstable understandings of how programs execute (sorva2013notional). Sentance and colleagues likewise emphasise the importance of explicit classroom talk about what code does, not only about writing it (sentance2023programming).
Program comprehension research has characterised code understanding as structural and compositional. Early work on procedural languages identified programming ”plans” - stereotyped code patterns that experts recognise and use to solve families of problems (soloway1984empirical), as summarised in later reviews (schulte2010block). Pennington’s studies showed that comprehending a program involves constructing both a control-flow oriented program model and a data-flow or situation model, and coordinating between them (pennington1987comprehension; schulte2010block). Schulte’s Block Model synthesises this work into a matrix with hierarchical levels (atoms, blocks, relations, macro) and dimensions of understanding (program text, execution, function) (schulte2008block). These frameworks articulate key abstractions that underpin expert comprehension, but say less about how meanings at different scales - from individual blocks to whole systems - are related in classroom discourse. In primary settings, teachers therefore still need practical ways of talking with pupils about structure, flow and purpose. Sentance and colleagues highlight that students and teachers require explicit language for explaining how programs work and why they behave as they do (sentance2023programming).
FLARE v1 (heath2025flare) responded to this need by adapting the Block Model for primary classrooms as four tiers for talking about block-based programs: Blocks, Segments, Relationships, and Macro. The CEP ’25 study showed that these tiers helped teachers structure questioning and classroom dialogue, but also exposed limitations. Teachers reported that the Relationships and Macro tiers were difficult to apply in simple programs and, at times, confusing; classroom analysis suggested that the Relationships tier in particular combined several distinct mechanisms, including temporal ordering, causal triggering, and shared state (heath2025flare). Introducing all four tiers at once also placed demands on teachers’ subject knowledge and working memory. This paper presents FLARE v2 as a recursive reformulation in which elements bind to form new elements at successive scales. The next section summarises the constraints that emerged in the CEP ’25 work before introducing the generative core of FLARE v2.
Conceptually, FLARE v2 can be read as an integrative extension of Pennington’s dual abstraction, Schulte’s Block Model, and Sorva’s account of notional machines in event-driven environments (pennington1987comprehension; schulte2008block; sorva2013notional). It aims to retain their insights while providing a compositional account of how program meanings are constructed across scales.
2 Legacy and Diagnosis: Constraints of FLARE v1
The CEP ’25 study demonstrated that FLARE v1 supported structured discussion but also revealed structural and pedagogical limitations when used in primary classrooms (heath2025flare). These limitations stemmed from how the fixed tier hierarchy interacted with teacher subject knowledge, artefact scale, and cognitive load.
2.1 Tier Granularity and Teacher Subject Knowledge
Teachers engaged readily with the lower tiers - Blocks and Segments - because these mapped onto visible code structures. Difficulty arose with Relationships and Macro, which required reasoning about invisible mechanisms such as triggers and shared state.
Teacher A stated: ”I didn’t understand half of them… There were elements that I was just totally baffled by.”
Teacher C similarly explained: ”Because I’ve not had any coding experience… you’re full, that’s you at capacity.”
These statements align with national evidence that many teachers lack specialist subject knowledge and feel under-supported in teaching the new curriculum (royalsociety2017, p. 72). The Royal Society’s 2025 review similarly highlights “uneven provision and lack of suitably qualified subject specific teachers” in computing education and notes that computing teacher recruitment remains critically below government targets (royalsociety2025). They also sit alongside findings that non-specialist teachers’ own subject knowledge and confidence shape how they interpret pupils’ programming difficulties (sentance2023programming).
2.2 Mismatch Between Tier Structure and Artefact Scale
FLARE v1 assumed that programs would exhibit meaningful structure at all four tiers. However, the block-based artefacts used in CEP ’25 - simple, event-driven scripts - rarely contained the long-range dependencies or global design features required for productive use of the Relationships and Macro tiers.
Teacher B noted: ”They weren’t linking into segments and therefore it wasn’t working… you’ve got two different segments that aren’t linked together.”
This pattern resonates with Schulte et al.’s distinction between ”programming in the large” and ”programming in the small”, where abstractions designed for large, complex systems may not apply straightforwardly to short instructional programs (schulte2010block, p. 75).
2.3 Cognitive Load and Simultaneous Introduction
Teachers reported that introducing all four tiers at once created cognitive load. Rather than functioning as interpretive lenses, the tiers sometimes became a checklist.
Teacher A reflected: ”Once I got going, it was like a snowball… but looking back at it, it was still overwhelming.”
Teacher C added: ”I didn’t have the capacity to develop the questions further… I didn’t have time to sit back and reflect.”
Classroom dialogue also showed that the Relationships tier conflated distinct mechanisms - temporal order, causal triggering, and state-sharing - leading to confusion when teachers attempted to apply it. In Black and Wiliam’s terms, teachers need time to invest in practices that deepen learning rather than simply pushing through content: they warn that ’”delivery” and ”coverage” with poor understanding are pointless and can even be harmful’ (black1998assessment, p. 144).
Table 1 summarises how the four FLARE v1 tiers map onto v2’s elements and bindings.
Table 1: Mapping FLARE v1 tiers to v2 elements and bindings
FLARE v1 Tier FLARE v2 Element FLARE v2 Binding
Blocks Elements at block scale (none)
Segments Elements at segment scale Causal-Temporal within segments
Relationships (none) Both binding dimensions
Macro Elements at system scale Purpose emerges from bindings
These limitations prompted a reconceptualisation. Where FLARE v1 was primarily a pedagogical scaffold - a set of tiers to organise classroom talk - FLARE v2 aims to be a generative theoretical account of how program meaning is constructed at any scale.
3 Formalising the Pattern: Code as Semiotic System
The limitations identified in FLARE v1 indicate the need for a descriptive lens that applies uniformly across scales, separates mechanisms rather than bundling them, and supports progressive introduction. Teachers’ natural reasoning about code focused on sequencing, triggering, conditions, and shared state, yet FLARE v1 did not articulate these distinctions.
FLARE v2 is designed to address these challenges by grounding the framework in a light semiotic footing. Classical structural linguistics emphasises that meaning in language arises from relations within a system and from the composition of bounded units into larger structures (saussure1916course; hjelmslev1961prolegomena). FLARE v2 adopts only a minimal commitment from this tradition: that bounded units carry meaning, and that binding these units produces new ones. This perspective supports the recursive element-binding cycle developed in the subsequent sections.
Figure 1 illustrates how literacy and programming share a common compositional structure. On the left, familiar literacy units (words, sentences, paragraphs, texts) combine through grammatical bindings. On the right, programming constructs (literals/expressions, commands, segments, systems) follow the same recursive pattern. This parallel provides teachers with an accessible metaphor: just as sentences combine to form paragraphs that express complete ideas, commands combine to form segments that accomplish discrete computational tasks.
Figure 1: The Compositional Ladder: parallel compositional progressions in literacy and programming. Both domains build meaning through recursive binding of smaller units into larger ones.
3.1 Design origins
The redesign of FLARE began from practice rather than theory. While using FLARE v1 with non-specialist teachers in block-based environments that sit on a JavaScript event loop, certain programs repeatedly exposed gaps in the framework. Simple scripts that mix non-blocking timers with sequential commands often behave in ways that pupils and teachers find surprising. For example, a program that schedules "morning" to be printed after one second and then immediately prints "Good" will output Good morning, even though many learners predict the opposite. FLARE v1 described structure at the level of Blocks, Segments, Relations and Macro, but it did not give teachers a precise way to talk about this kind of timing behaviour, about non-blocking constructs, or about the event loop that underpins them. The Relations tier in particular blurred temporal order, causal triggering and shared state, and classroom transcripts showed the same mechanisms being invoked in both the Segments and Relations tiers.
In response, I set aside high-level classroom examples and worked downwards to smaller and smaller representations, asking what the most fundamental units of program structure might be. This led to a deliberately reductive exercise with a small RISC-style instruction set. Registers and memory locations were treated as ”things”, core instructions such as MOV, ADD, SUB, JMP and CMP as actions on those things, and branch instructions such as BEQ and BNE as choices after a comparison. The immediate goal was not to recover familiar control structures, but to see whether this reduction could separate the conflated concerns in FLARE v1, particularly the mix of temporal sequencing and logical control that appeared in both the Segments and Relations tiers.
What emerged was a fundamental distinction in how control moves between instructions. Code that executes without any branch instruction - where the instruction pointer simply increments - differs categorically from code that requires a jump to a label. Sequential flow needs no branch; conditionals, loops, and function calls all require one. This binary provided a cleaner foundation than FLARE v1’s conflated Relations tier.
However, a further distinction became apparent when considering event-driven code. When a child writes an event handler such as button.onClick(...), the handler is registered, but invocation is controlled by the runtime’s event loop - an invisible while(true) that checks for events and dispatches handlers. Mechanically, this is just another branch inside a loop. But the child did not write that loop; they cannot see it. From the learner’s perspective, the handler ”just fires” when clicked. This suggested that event-driven control, while mechanically equivalent to branching, deserves separate treatment because the branching mechanism is hidden from the programmer.
A subsequent exploratory classroom probe (n=26, ages 9-10) revealed an interesting pattern. When shown event-driven Scratch-style code and asked what would happen if a second event fired mid-execution, 77% of pupils selected concurrent execution on the probe items, without prompts or worked examples in the session itself. These children had most recently encountered an event-handling metaphor two years earlier in Year 3, when we talked about indented code as ”levels” and they proposed an event-handling metaphor, noting that multiple handlers could work simultaneously. They had received no further teaching on concurrency in the intervening years. This suggests that such early metaphors may form lasting execution models, and that children in event-driven, block-based curricula can reason concurrently about event-driven code, even though the classical comprehension studies on which this work builds, such as Pennington’s FORTRAN studies, focused on sequential control-flow abstractions (pennington1987comprehension). Full methodology and results are reported in heath2025concurrency.
Pennington’s work on procedural FORTRAN documented how programmers in single-threaded languages reason about control-flow and data-flow within a sequential execution model (pennington1987comprehension). The finding that children reason concurrently without explicit instruction suggests that this empirical regularity may not generalise straightforwardly to learners whose first programming experiences are event-driven.
Two design commitments for FLARE v2 followed from this work. First, any construct that consistently behaves as a single unit in reasoning or classroom dialogue should be treated as an element, provided it can be described in terms of what it receives, sends, effects and shares. Second, the ways elements connect should be separated into two orthogonal dimensions: Causal-Temporal binding (how control moves, with Sequential, Branch, and Event subtypes) and Communicative binding (how data flows). Only after this design work did I encounter Pennington’s distinction between control-flow and data-flow representations (pennington1987comprehension). The alignment between that work and the binding scheme described here is therefore a convergence that helps situate FLARE v2 within existing theory, rather than the origin of the framework itself.
4 The Generative Core of FLARE v2
The Compositional Ladder and framing developed in Sections 2 and 3 suggest a simple generative move that recurs across scales: elements bind to form new elements, which can themselves be treated as units in further bindings. Figure 2 visualises this recursive cycle, and FLARE v2 formalises this move in two parts:
  1. 1.
    a description of how individual elements can be expressed through four observable properties; and
  2. 2.
    a set of binding dimensions that explain how elements combine to yield new bounded units.
Together these form a recursive element-binding-element cycle that applies to blocks, segments, systems, and larger structures.
Figure 2: The generative core of FLARE v2: elements at scale $n$ bind to form new elements at scale $n+1$. The same descriptive properties and binding mechanisms apply recursively at every level.
4.1 Elements and their four properties
FLARE v2 uses the term element for any bounded unit of code that can be treated as a single meaningful entity in classroom discussion. As summarised in Table 1, constructs that appeared as separate tiers in FLARE v1 (for example, Blocks and Segments) are treated in v2 as elements at different scales but described using the same four properties.
Depending on context, an element may be a block, statement, function, script, subsystem, or larger component. At every scale, the same four observable properties apply:
  • Receives: inputs or triggers that cross the element boundary, such as events, parameters, sensor readings, inbound values, or user inputs.
  • Sends: outputs that cross the boundary, such as return values, emitted events, messages, or hardware commands.
  • Effects: internal state changes confined within the element, such as updates to local variables or private data structures.
  • Shares: any cross-boundary state or resource that the element reads or writes, such as global variables, configuration values, device state, shared hardware pins, or external services.
The distinction between Effects and Shares is crucial. Effects are private and unobservable by other elements. Any change that can be observed outside the element - because it modifies shared state or interacts with hardware - is a Share. Effects describe private state; Shares describe public or external state. Transient messages or one-off calls that pass values between elements are treated as Sends, not Shares. Shares are reserved for persistent cross-boundary state or resources that can be read or written by multiple elements over time.
An important consequence of this distinction is that Effects do not participate in bindings between elements. Because Effects are private and unobservable, they cannot create relationships with other elements. Only Receives, Sends, and Shares generate the inter-element connections that constitute Communicative bindings. Effects matter for understanding what happens inside an element, but they do not contribute to how elements combine.
Table 2 illustrates this distinction across different scales.
Table 2: Effects and Shares across scales
Scale Effects (private state) Shares (cross-boundary state)
Single block Updates a local counter used only inside the handler. Writes to a global score variable or hardware pin.
Segment (function / handler) Manipulates a private list or calculation scratch space accessible only within the segment. Reads or writes a global threshold value or updates a device such as a microcontroller LED grid.
System Maintains an internal scheduling index or mode flag used only within the system. Interacts with shared configuration, logs to a cloud service, or changes hardware output.
Because the same four properties apply at every level, teachers can move seamlessly between discussing a single block and a whole program. In classroom use, these properties become question stems: ”What does this element receive?”, ”What does it send?”, ”What does it change inside itself?”, and ”What does it share?” The vocabulary remains the same regardless of scale.
Figure 3 shows how FLARE v2’s recursive structure can map across different programming paradigms and levels of abstraction. The framework applies equally to visual block-based languages, text-based languages, and even lower-level representations like assembly code. I argue that this apparent universality emerges because the element-binding pattern is a fundamental way to describe how computational meaning is constructed, regardless of surface syntax, at least in the imperative and event-driven contexts typical of school curricula.
Figure 3: FLARE v2 structural layers showing how elements exist at multiple scales, from sub-blocks through to ecosystems. The same recursive grammar applies at each level.
4.2 Binding dimensions: how elements combine
Section 2 identified that FLARE v1 conflated several distinct mechanisms under a single Relations tier. The design work described in Section 3 revealed that these mechanisms separate cleanly into two orthogonal dimensions, closely aligned with Pennington’s distinction between control-flow and data-flow abstractions (pennington1987comprehension).
FLARE v2 retains Schulte’s Block Model insight that comprehension involves structure, flow, and function (schulte2008block), but refactors these concerns into element properties (Receives, Sends, Effects, Shares) and binding dimensions (Causal-Temporal and Communicative). Two additions distinguish the present framework: recursion, which the Block Model does not formalise; and explicit treatment of Event binding, which accommodates event-driven paradigms that the Block Model’s imperative orientation does not directly address.
The dimensions are orthogonal in the sense that for any pair of elements, two independent questions can be asked: how does control move between them (Causal-Temporal), and how does information move (Communicative). The answers to these questions can vary independently.
4.2.1 Causal-Temporal binding: how control moves
The Causal-Temporal dimension describes how execution flows between elements. The assembly-level analysis in Section 3 revealed three distinct subtypes based on the underlying control mechanism:
  • Sequential: No branch instruction is required; the instruction pointer simply increments. One element completes, the next begins. This is the default flow within a segment when no conditional or loop intervenes.
  • Branch: A jump instruction exists in the user’s code - whether for conditionals (if/else), loops (while, repeat), or function calls. The programmer wrote the control structure, and it is visible in their code.
  • Event: The jump exists, but it is hidden inside the runtime’s event loop. When a child writes button.onClick(handler), they write only the handler; the dispatching loop that decides when to invoke it is invisible. Mechanically, Event reduces to Branch inside a loop - but from the learner’s perspective, the handler ”just fires” in response to external stimuli. The classification here is from the learner’s perspective: the hidden loop is precisely why Event warrants separate pedagogical treatment, even though it reduces mechanically to Branch.
The exploratory Year 5 probe reported in Section 3 found that 77% of pupils selected concurrent execution when a second event was triggered. These pupils had encountered an event-handling metaphor two years earlier in Year 3, and noted multiple handlers working at the same time, but they had received no further teaching on concurrency in the intervening years. This suggests that such early metaphors may support concurrent execution models that persist and transfer, rather than emerging only from later formal instruction.
4.2.2 Communicative binding: how data moves
The Communicative dimension describes how information flows between elements. It operates through three of the four element properties: Sends, Receives, and Shares. (Effects, being private, do not participate in inter-element binding.)
  • Sends/Receives: Transient data flow - return values, parameters, messages passed between elements.
  • Shares: Persistent state accessible to multiple elements - global variables, hardware state, configuration values.
Where the original FLARE v1 Relations tier bundled ”what triggers this” with ”what does this share”, the two-dimension scheme separates them cleanly. Causal-Temporal binding answers when and why control moves; Communicative binding answers what information moves and where it persists.
4.2.3 Why two dimensions, not three?
An earlier formulation of FLARE v2 proposed three binding dimensions: Causal-Temporal, Logical-Conditional, and Communicative-Referential. The Logical-Conditional dimension was intended to capture predicates, guards, and Boolean conditions. However, further analysis revealed that conditionals are not a separate kind of binding but rather a mechanism that implements branching control flow. The if block, the loop guard, and the Boolean expression are all constructs that produce Branch-subtype Causal-Temporal bindings.
Treating conditionals as a separate dimension conflated the mechanism (Boolean evaluation) with the effect (control transfer). The revised scheme places conditionals where they belong: as the means by which Branch bindings are created. This simplifies the framework without losing descriptive power. Questions about conditions (”Under what circumstances does this run?”) remain important but are now understood as questions about how a Branch binding is governed, not as a separate binding dimension.
Table 3 summarises the two dimensions, their subtypes, and representative question stems.
Table 3: Binding dimensions: subtypes, mechanisms, and example question stems
Dimension Subtype Mechanism Example question stems
Causal-Temporal Sequential No branch; instruction pointer simply increments. What happens next?
What runs after this?
Branch Jump in user code (conditionals, loops, calls). Under what conditions does this run?
How many times?
Event Jump hidden in runtime loop. What triggers this?
When does this fire?
Communicative (none) Data exchange via Sends, Receives, Shares. What does this receive?
What does it send?
What do these share?
These dimensions are not proposed as an exhaustive taxonomy; other patterns may matter in object-oriented, concurrent, or distributed systems. But these two dimensions with three Causal-Temporal subtypes have proved sufficient in my own practice for supporting reasoning and classroom talk in imperative and event-driven contexts typical of primary and lower-secondary classrooms.
4.3 The recursive element-binding-element cycle
With elements and binding dimensions established, the core of FLARE v2 reduces to a four-step recursive cycle:
  1. 1.
    Identify elements at the current scale. Decide what counts as an element for the current question.
  2. 2.
    Describe each element using the four properties. Ask what it receives, sends, effects, and shares.
  3. 3.
    Analyse bindings between the elements. Examine Causal-Temporal connections (Sequential, Branch, or Event) and Communicative connections (what flows between them, what they share).
  4. 4.
    Recognise the new element that emerges. Treat the bound set as a new element at the next level, with its own properties, capable of further binding.
This cycle can be applied repeatedly. At one level, it explains how blocks bind into a segment; at the next, how segments bind into a system. Nothing in the cycle depends on visual form: it applies equally to block-based, text-based, and hybrid environments.
In classroom practice, the cycle is enacted through question stems rather than formal steps. A teacher might first stabilise the elements in view (”Which bits belong together?”) before moving to questions about what each receives or sends, then to when and under what conditions interactions occur, and finally to naming the new unit that emerges. Section 5 illustrates this on a concrete program.
It is important to distinguish FLARE v2’s formal grammar from its classroom realisation. The analytic labels - Causal-Temporal, Communicative, Sequential, Branch, Event - are not intended as classroom vocabulary. In practice, these appear as simple, scalable questions such as ”What happens next?”, ”What triggers this?”, ”Under what conditions?”, and ”What does this share?” The cognitive load reported in FLARE v1 arose partly from presenting a multi-tier structure all at once. FLARE v2 supports progressive reveal: the same family of questions can be rehearsed first on blocks, then on segments, and then on systems, without introducing new structural categories at each stage.
4.4 Relationship to existing program comprehension models
FLARE v2 can be situated relative to three established traditions in program comprehension research. It aligns with Pennington’s program-versus-situation model, Schulte’s Block Model, and Sorva’s account of notional machines, while extending each to event-driven environments typical of contemporary school curricula.
Pennington’s dual abstraction.
Pennington showed that programmers construct two complementary mental representations: a program model that foregrounds control flow, and a situation model that foregrounds data flow and goals (pennington1987comprehension). FLARE v2 mirrors this duality in its two binding dimensions: Causal-Temporal binding captures how control moves, while Communicative binding captures how information and shared state move between elements. The main extension is to make event-driven control an explicit subtype, acknowledging hidden runtime mechanisms that Pennington’s procedural studies did not need to model.
Schulte’s Block Model.
The Block Model conceptualises program comprehension as reasoning about multiple hierarchical levels (from atoms to macrostructure) along three dimensions: text surface, execution, and function (schulte2008block). FLARE v2 preserves this insight but refactors it: structure and flow are captured by the two binding dimensions, while function (purpose) is treated as an interpretive stance that can be applied at any scale. The key addition is the recursive element-binding-element cycle, which explains how meaning propagates between levels rather than treating levels as fixed.
Sorva’s notional machines.
Sorva reviews evidence that learners form mental models of program execution whether or not explicit notional machines are taught, and that hidden runtime mechanisms are a major source of misconception in many environments (sorva2013notional). FLARE v2 responds by making Event binding explicit as part of the Causal-Temporal dimension and by separating visible and hidden control mechanisms. This gives teachers a way to surface the invisible event loops and schedulers that govern much of the block-based code children encounter, while remaining agnostic about the particular notional machine adopted.
Taken together, these correspondences position FLARE v2 as an integrative extension of classical procedural comprehension models to the event-driven, often concurrent artefacts used in school computing.
5 Illustrative Example: A Simple Control System
This section illustrates the element-binding-element cycle and the four properties using a minimal three-segment control system commonly encountered in introductory physical computing and robotics. The system, which we refer to as AutoDimmer, comprises a ReadSensor segment that samples the current light level, a DecideBrightness segment that determines an appropriate LED brightness based on a threshold, and a SetLED segment that applies the chosen brightness to the hardware. This sense-decide-act pattern is familiar in many classroom microcontroller projects. The example is intentionally simple and language-agnostic: its purpose is to make the element-binding-element cycle concrete rather than to present a complex or novel control design.
Figure 4 shows the three segments identified as bounded elements, with their bindings forming the AutoDimmer control system.
Figure 4: Three segments identified as bounded elements. Their bindings form a simple
AutoDimmer control system, illustrated here before the formal analysis in Table 4.
5.1 Step 1: Identify elements at the current scale
On the first pass through the FLARE cycle, the elements in view are the three segments: ReadSensor, DecideBrightness, and SetLED. In a classroom, a teacher might stabilise these by asking ”Which bits of code belong together here?” and then marking or naming each segment as a unit.
5.2 Step 2: Describe each segment with the four properties
Each segment can then be described using Receives, Sends, Effects, and Shares. Table 4 summarises this for the three segments.
At this stage, a teacher’s questions might be:
  • ”What does ReadSensor receive? What does it send? Does it change anything inside itself?”
  • ”What does DecideBrightness depend on? What shared values does it read?”
  • ”What does SetLED change in the world? Does anything else share that device?”
These questions apply the four properties to each segment without yet considering how the segments are connected.
Table 4: Four properties for each segment in the control system
Segment Receives Sends Effects Shares
ReadSensor Trigger from loop or timer Numeric light level None Reads sensor hardware state
DecideBrightness Light level value Chosen brightness value Temporary internal variables only Reads the shared threshold value
SetLED Brightness value Command to LED hardware None (no private state) Writes to LED device (shared hardware)
5.3 Step 3: Analyse bindings between segments
The next step is to analyse how the segments bind together along the two binding dimensions. The same three segments illustrate both dimensions and multiple Causal-Temporal subtypes.
Causal-Temporal binding.
The three segments exhibit different Causal-Temporal subtypes:
  • Sequential binding connects the segments within each cycle: ReadSensor completes, then DecideBrightness runs, then SetLED executes. No branch instruction governs these transitions; each segment simply follows the previous one.
  • Branch binding operates within DecideBrightness, where a conditional compares the sensor reading with the threshold and selects between alternative brightness outputs. The if/else construct creates a branch in the user’s visible code.
  • Event binding (or a timer-based equivalent) governs when the entire sense-decide-act cycle begins. If a forever loop or timer tick triggers each cycle, this trigger mechanism is often partially hidden - the learner writes ”forever” but may not see the underlying loop mechanics. In a purely event-driven version, a sensor-change event might trigger the cycle directly.
Typical questions for this dimension include:
  • ”What has to happen before DecideBrightness can do its job?” (Sequential)
  • ”Under what conditions does the LED get set to maximum?” (Branch)
  • ”What triggers this whole chain to start each time?” (Event or timer)
Communicative binding.
The elements are linked by the values they exchange and the state they share:
  • ReadSensor sends the sampled light level; DecideBrightness receives it.
  • DecideBrightness sends a chosen brightness value; SetLED receives it.
  • The threshold value that DecideBrightness reads is a shared configuration that other elements in the wider program could also access or update.
  • The LED device that SetLED writes to is shared hardware - other subsystems might also control it.
Questions for this dimension include:
  • ”Which segments rely on, or pass along, the light level value?”
  • ”Where does the threshold come from, and who else can change it?”
  • ”What would happen if another part of the program also tried to control the LED?”
Note that the internal calculations within DecideBrightness - any temporary variables used to compute the brightness-are Effects and do not participate in bindings. They matter for understanding what happens inside the segment, but they create no relationships with other elements.
Taken together, these bindings show why the three segments form a coherent control subsystem rather than three unrelated actions. Both binding dimensions operate concurrently. The same conditional that creates a Branch binding (comparing light level to threshold) also depends on Communicative binding (reading the shared threshold and the received light level). In classroom talk, a teacher might foreground one dimension at a time, typically starting with Sequential flow, then layering in questions about conditions (Branch) and shared state (Communicative).
5.4 Step 4: Recognise the system as a new element
Once the bindings are understood, the three segments can be treated as a single element at the system layer: the AutoDimmer system. From this point on, the system can be described using the same four properties:
  • Receives: ambient light readings (from the sensor) and periodic triggers from the control loop or timer.
  • Sends: brightness commands addressed to the LED hardware.
  • Effects: optional internal bookkeeping such as smoothing counters or mode flags. In this minimal example, the system has no private internal state.
  • Shares: external resources it interacts with, including the LED device and the shared threshold used by the decision logic.
The physical change in LED brightness is not an internal Effect of the AutoDimmer system. It is a change to an external device and therefore a Share. The system sends a brightness command; the resulting change in device state is a write to shared external state.
Recognising the system as a new element completes one turn of the recursive cycle: blocks combine into segments; segments combine into a system; and the resulting system can now participate in further bindings at a higher scale.
5.5 Stepping up the ladder: system within a larger service
The same cycle applies at higher scales. In a typical microcontroller project, AutoDimmer might sit within a larger room-controller service alongside subsystems such as HeatControl and PresenceDetect. At this scale, elements are whole systems rather than individual segments.
A room-controller service might:
  • Receive sensor data from light, temperature, and motion detectors, plus external commands from a user interface or network.
  • Send commands to lighting, heating, and display devices, or outbound messages to a cloud service.
  • Effect internal scheduling or mode flags that are private to the service.
  • Share configuration, occupancy state, or hardware devices that other subsystems may also read or update.
The bindings between subsystems follow the same two dimensions:
  • Causal-Temporal: Which subsystem runs first? What events trigger each? Do any subsystems wait for others to complete (Sequential), respond to conditions set by others (Branch), or activate independently on external stimuli (Event)?
  • Communicative: What data flows between subsystems? What configuration or hardware state do they share?
Changes to heaters, LEDs, displays, or network endpoints are all writes to shared state, not internal Effects. The definitions of Receives, Sends, Effects, and Shares therefore remain stable as the recursion climbs the ladder.
6 Discussion: Pedagogical and Curricular Implications
The Compositional Ladder suggests a natural pattern of progression: learners first gain fluency with smaller elements and simple bindings, then encounter larger elements whose internal structure can be analysed using the same questions. This aligns with established principles of spiral curricula in which ideas are revisited with increasing sophistication. Rather than introducing a full hierarchy at once, teachers can dwell at a particular rung until learners are fluent with its elements and bindings, and then move upward when appropriate.
FLARE may be understood both as a descriptive grammar for analysing program structure and as a mental framework that may shape how teachers scaffold the construction of code. By providing a stable set of properties (Receives, Sends, Effects, Shares) and binding dimensions (Causal-Temporal with its three subtypes, and Communicative), FLARE offers novices a way to articulate what a segment is intended to do before or while constructing it. To be clear: FLARE v2 is not a psychological model of novice cognition. It is a semiotic-compositional account of program meaning intended to structure teacher questioning and learner reasoning.
The separation of Causal-Temporal subtypes may prove particularly valuable for event-driven environments. The Year 5 probe reported in Section 3 found that most children selected concurrent execution when a second event was triggered. Given that their only prior exposure had been a brief event-handling metaphor two years earlier, with no reinforcement since, this pattern suggests that early metaphors can support persistent concurrent execution models in event-driven environments.
Because the ladder parallels familiar literacy progressions, FLARE may reduce cognitive load for teachers whose professional expertise lies in literacy, though this remains to be tested empirically.
From a pedagogical perspective, PRIMM (sentance2019primm) offers a temporal sequence for classroom activity (Predict, Run, Investigate, Modify, Make). FLARE may complement this. PRIMM structures what learners do over time; FLARE provides a way to talk about what they see and think at each phase.
A simple illustration shows how the two can interact. During Predict, a teacher might ask pupils which elements they expect to be involved in a behaviour and what triggers them (”Which segments will run when the button is clicked?”-an Event binding question). In Run, these elements are observed and named. During Investigate, FLARE questions focus attention on Causal-Temporal flow (”What happens next? What conditions control this branch?”) and Communicative connections (”What does this segment share?”). In Modify and Make, pupils reuse the same vocabulary to assemble known elements into new arrangements (”What will this new segment receive and send?”). On this view, PRIMM provides the classroom rhythm, while FLARE supplies the underlying grammar that helps that rhythm culminate in genuinely compositional, creative work. These relationships remain speculative and require empirical evaluation.
In curriculum design, the recursive model suggests a spiral structure in which new abstraction levels appear once earlier bindings are secure. Early years may emphasise Sequential flow within small elements; later years can introduce Branch bindings with their associated conditional reasoning, and then Event bindings with their hidden control mechanisms. This mirrors literacy development and may support teachers in extending familiar practices from reading and writing text to reading and ”writing” code structure. Because Macro is treated as an interpretive stance rather than a separate tier, questions of purpose and audience (”What is this element for?”, ”Who relies on it?”) can be introduced at any rung without adding another structural category.
7 Limitations and Future Work
This paper is conceptual and exploratory. It does not report new empirical data beyond the exploratory Year 5 probe reported in Section 3 and detailed in heath2025concurrency. That probe requires systematic replication with controlled conditions. The analysis is interpretive and practitioner-led, drawing on reflective re-reading of earlier classroom work rather than on a newly collected or systematically coded dataset.
The two-dimension binding scheme with three Causal-Temporal subtypes has been developed for imperative and event-driven contexts typical of school curricula. Its applicability to object-oriented, concurrent, or distributed paradigms remains untested. The claim that Event binding is ”mechanically equivalent to Branch but pedagogically distinct” requires further investigation: do learners actually benefit from this distinction, or does it introduce unnecessary complexity?
Future research should examine how the recursive binding cycle manifests in teacher and learner discourse and whether it transfers across languages, environments, and programming paradigms. Studies might analyse classroom dialogue using the element–binding–element cycle and the two binding dimensions, investigate how teachers use the Compositional Ladder to structure tasks and questions, or explore whether making Event binding explicit helps learners develop more accurate notional machines for event-driven code. Implementation studies and evaluations of classroom uptake lie beyond the present scope but will be essential for assessing the practical value of FLARE v2.
8 Conclusion
The shift from a four-tier descriptive hierarchy (Blocks, Segments, Relations, Macro) to a recursive element–binding cycle represents a move from classification toward generative explanation. FLARE v2 positions program comprehension as a semiotic process: bounded elements carry meaning, and new meaning emerges as elements bind through Causal-Temporal relations (Sequential, Branch, or Event) and Communicative relations (what flows between elements and what they share). This reframing provides a unified conceptual vocabulary for discussing structure, flow, and purpose across artefact scales.
The two-dimension scheme aligns with Pennington’s established distinction between control-flow and data-flow abstractions. Although Pennington’s original work concerned sequential procedural programs, the control-flow / data-flow distinction can be generalised to event-driven environments where control mechanisms are hidden from the learner. By treating Event as a distinct subtype of Causal-Temporal binding, FLARE v2 acknowledges both the mechanical reality (events are dispatched by branches in a runtime loop) and the pedagogical reality (learners cannot see that loop and may develop intuitions that differ from sequential-first assumptions).
FLARE v2 emerged from reflective classroom experience and is presented here as a theoretical proposal rather than an empirical claim. In computing education, practitioner-informed conceptual work serves an important role: it can articulate implicit classroom reasoning, connect practice with broader theoretical traditions, and motivate targeted empirical studies. The framework introduced here invites such investigation. Future work should examine how the element–binding cycle can support teacher questioning, pupil reasoning, and code creation; how it complements existing pedagogical approaches such as PRIMM; and how it may inform curriculum design across primary and secondary phases.
In this sense, the purpose of FLARE v2 is not to settle questions of program comprehension, but to open them. By offering a stable, scalable vocabulary grounded in both semiotic theory and classroom need, it aims to provide a basis for subsequent empirical evaluation, refinement, and dialogue within the computing education community.
9 Acknowledgements
With thanks to Robert Whyte and Sue Sentance, whose guidance and collaboration during the TICE2 project and CEP ’25 study laid the foundations for the ongoing development of FLARE.
10 Funding
The author participates in the Microsoft for Startups Founders Hub programme, which provided Azure credits used for technical prototyping of a tool that applies the FLARE framework. No external funding supported the research reported in this paper.
11 Conflict of Interest
The author is developing a tool that applies the FLARE framework, which may be commercialised in the future. The work reported in this paper was completed independently of the tool’s development. No other conflicts of interest are declared.
Licence
This preprint is made available under the Creative Commons Attribution–NonCommercial–ShareAlike 4.0 International Licence (CC BY-NC-SA 4.0). This permits copying, distribution, and adaptation for non-commercial purposes, provided that appropriate credit is given and that derivative works are released under the same licence.
This licence applies only to the text, figures, and explanations contained in this paper. Practical implementations of the framework—including training materials, classroom scaffolds, exemplar resources, and software tools—are licensed separately and are not covered by this CC BY-NC-SA 4.0 licence.
Author’s intention.
FLARE is released openly to support teachers’ professional judgement rather than prescribe scripted schemes of work. The intention is to strengthen educator autonomy, not to constrain it.